OpenStudy (anonymous):

how you evaluate the expression? 9!/7! * 2! is it 36?

OpenStudy (anonymous):

\[\frac{ 9! }{ 7!*2 } = \frac{ 9*8*7! }{ 7!*2 } = 36\]

OpenStudy (anonymous):

thanks!